40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Aug 24 18:37:37 2010, Zach, Laser, GYRO, Gyro "single-arm" test single_arm.png
    Reply  Tue Aug 24 19:24:58 2010, rana, Laser, GYRO, Gyro "single-arm" test 
       Reply  Tue Aug 24 20:06:16 2010, Zach, Laser, GYRO, Gyro "single-arm" test 
          Reply  Wed Aug 25 01:02:17 2010, rana, Laser, GYRO, Gyro "single-arm" test 
          Reply  Wed Aug 25 01:09:43 2010, Zach, Laser, GYRO, Gyro "single-arm" test single_leg_noise.png
Message ID: 976     Entry time: Tue Aug 24 18:37:37 2010     Reply to this: 978
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: Gyro "single-arm" test 

 After yesterday's measurement, it dawned on me that the noise measured in the free-running MZ setup is not the only kind of displacement noise we are worried about. It tells us something about the high-frequency shaking of the optics but does not tell us anything about the lower-frequency "breathing mode" type behavior as this is common to both MZ legs. While the gyro is still in pieces (sorry Alastair), I thought I might run a single-arm test, gyro style.

Simply put, I made a linear cavity out of one leg of the gyro, using the big input mirror and the southwestern curved mirror (the one by the laser). This is easy to do because the gyro modematching solution also works for this setup. The idea here is to let it run for some time and measure the low-frequency length drift. Of course, the assumption is that the expansion of the table is isotropic, so that this length drift is an indication of how the area of the cavity will change over long timescales. I don't see why this should not be the case.

Below is a photo diagram of the setup. I left the CW path blocked upstream of its faraday isolator, then installed a PBS/QWP isolator before the last steering mirror into the cavity on the CCW path. The QWP was stolen from the AOM double-pass setup temporarily. I was able to set up the REFL optics in such a way that the normal gyro optics are unaffected; with any luck, we will need no additional realignment as a result of this test.


The REFL PD is a PDA10A, and the modulation frequency is 33 MHz. I have the slow and fast controls signals acquired, as well as the transmission, so that I can monitor it overnight. I can force the thing to re-lock by messing with the offset to the slow control. This should not be necessary, though, as for some reason we have a MUCH stabler lock with the linear cavity. I have not taken any transfer functions, but it seems clear that we have much higher bandwidth, as I can knock on the table--hard--with no problems. Of course, the cavity is now highly overcoupled, which I believe gives us more optical gain, but in this configuration I can turn the PDH box gain way up before seeing any resonances. The fact that I'm using a 150-MHz REFL PD probably doesn't hurt. This might be reason to believe that our crappy lock issue is PD-related.

ELOG V3.1.3-